# Inverse Transition Learning: Learning Dynamics from Demonstrations

### v1

@@ Introduction
Environments Our evaluations span 11 synthetic environments and a real-life ICU setting. Specifically, we use a 25 states Gridworld with four actions, 10 different 15-states Randomworld with five actions each, and a healthcare scenario focusing on ICU patients with hypotension, utilizing the MIMIC-IV dataset (Johnson et al. 2020). In Gridworld, we generate 100 batches of data $\mathcal{D}$, each consisting of five episodes with 15 steps each. For each 10 of the Randomworld, 50 batches are generated, each containing three episodes of ten steps, due to their smaller size. We also evaluate our methods across varying ”Coverage %” levels in these environments, defined as the percentage of states observed within each batch data $\mathcal{D}$ and various $\epsilon$ values to see how our method compare to baselines for various degrees of sub-optimality. Detailed descriptions of both the synthetic and real-world environments are available in the Appendix.
@@
In synthetic scenarios, we simulate a suboptimal expert, where approximately $40\%$ of states are stochastic-policy states (See definition 4), demonstrating the robustness of our approach to suboptimal expert behavior. We also include in the Appendix (Table 5, 6, Figure 6 and 7), results for $20\%$ and $0\%$ (fully optimal expert) of stochastic-policy states which correspond to lower $\epsilon$ values. Note that increased optimality from the expert (lower $\epsilon$ values) leads to less stochasticity (ie. less stochastic-policy states) amongst the actions selected in the batch data $\mathcal{D}$. Our method does not require the expert to be perfectly optimal and can adapt to various degrees of optimality through $\epsilon$. In the results below, we use $\gamma=0.95$ and smoothing parameter defined in Equation 5 $\delta=0.001$.
@@
Baselines. Our methods, Bayesian Inverse Transition Learning (BITL) and Inverse Transition Learning (ITL), are evaluated against $T^{MLE}$ (Maximum Likelihood Estimation, MLE), $T^{MCE}$ (Maximum Causal Entropy, MCE) (Herman et al. 2016), and a non-expert-informed posterior $P(T^{*}|\mathcal{D})$ (Posterior Sampling, PS).
@@
Best/$\epsilon$-ball action matching: Proportion of states where the best/$\epsilon$-good action taken:
$\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}\mathcal{I}\{\arg\max_{a}\pi(a|s)=\arg\max_{a}\pi^{*}(a|s)\}\Big{/}\\ \frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}\mathcal{I}\{\arg\max_{a}\pi(a|s)\in\epsilon(s;T^{*})\}$
@@
$\frac{1}{|\widehat{P}(T|\mathcal{D})|^{2}}\sum_{T\in\widehat{P}(T|\mathcal{D})}\sum_{T^{\prime}\in\widehat{P}(T|\mathcal{D})}\\ \mathbb{E}_{s_{0}\sim\mu_{0}}\left[|V^{\pi^{*}(T)}(s_{0};T)-V^{\pi^{*}(T^{\prime})}(s_{0};T)|\right]$
We also report the normalized Value, calculated as the ratio of the value achieved by the optimal policy of our learned dynamics under the true dynamics $T^{*}$ to the expected value under the true dynamics $\mathbb{E}_{s_{0}}[V^{*}(s_{0};T^{*})]$. Finally, we report the CVaR of Value over datasets, Total Variation, the number of our constraints (8, 9) violated, and the training Time(in seconds) for the baseline MCE and our method ITL. Value and Best/$\epsilon$-ball action matching metrics are computed both in standard and transfer settings. In this paper, a transfer task refers to the evaluation of the learned dynamics and/or policy under a reward function different from the one used during training. To demonstrate that our learned dynamics can adapt to actions not seen in the demonstrations, we set up a transfer task where the optimal actions under this new reward function differ from those in the original task (See Figure 5, 6). This ensures the new task requires actions we didn’t observe in the initial demonstrations. This tests the model’s adaptability to changes in task conditions.
Computing the results. For each dynamics $T$, we compute its optimal policy $\pi^{*}(T)$ using Value Iteration (Sutton and Barto 2018), when needed to compute a metrics in terms of policies. For posterior distributions $P(T|\mathcal{D})$ and $P_{\epsilon}(T|\mathcal{D})$, we determine the optimal policies of 5,000 sampled $T^{\prime}s$ and average them.
Figure 1: Performance of ITL on a held out validation set across different $\epsilon$ values.
How to Tune $\epsilon$: In the healthcare setting, we only have access to an offline dataset. To select an appropriate $\epsilon$, we evaluate the performance of ITL on a held-out validation set. We choose the $\epsilon$ that performs best on this validation set based on the Best/$\epsilon$-ball action matching metric for both standard and transfer tasks (defined by different reward functions). As shown in Figure 1, $\epsilon=5$ appears to be suitable, and this value is used for training and computing the results in the healthcare experiments (See Table 1). We also present results for $\epsilon=10$ and $\epsilon=15$, as these values are also reasonable. Results for $\epsilon=10$ and $\epsilon=15$ are provided in Tables 7 and 8 in the Appendix, showing similar performance.
Results
Figure 2: Top row: Normalized Value vs. Coverage for Gridworld (left: Standard Task, middle: Transfer Task), Bottom row: Normalized Value vs. Coverage for Randomworlds (left: Standard Task, middle: Transfer Task). Rightmost plots: Normalized Value vs. Bayesian Regret of both Tasks (top: Gridworld, bottom: Randomworlds).
Figure 3: Most likely next 3 states after prescribing Intravenous treatment in state O2= 1, BP = 1, GCS = 1, Crea = 2
Table 1: Healthcare dataset results environment ($\epsilon=5$)
Standard Task Transfer Task
Method MLE ITL MCE BITL PS MLE ITL MCE BITL PS
Best matching 0.33 0.51 0.31 0.49 0.31 0.34 0.52 0.32 0.47 0.34
$\epsilon$ matching 0.52 1 0.56 1 0.51 0.58 0.97 0.68 0.90 0.58
Nbr Constraints 61 0 47 0 58 - - - - -
Time - 2.23 118 - - - - - - -
Bayesian Regret - - - 2.49 10 - - - 2.80 5.40
ITL consistently outperforms baseline methods across all metrics and coverage settings. Table 4, Table 1 and Figure 2 show that ITL and BITL outperform other baselines across different metrics and coverage levels, in both synthetic environments and the real-life healthcare environment. In Randomworlds, BITL and ITL significantly surpass MCE, which outperforms MLE and PS. In Gridworld, ITL slightly outperforms MCE on average but exhibits significantly better worst-case performance. This highlights the importance of inferring dynamics that satisfy hard constraints and the potential of MCE to get stuck in bad local optima due to gradient-based optimization. ITL also takes significantly less time to train. BITL outperforms all methods, showing that modeling uncertainty over dynamics while leveraging expert demonstrations can help outperform point estimate methods. In the standard task setting, as Theorem 1 guarantees, ITL and BITL are the only methods converging to optimal as coverage increases (left column of Figure 2). They achieve higher values without violating constraints, attributed to enforcing hard constraints in our work versus soft constraints in MCE.
ITL enables reliable, fast, and data-efficient dynamics estimation. In Gridworld, MCE required significantly more training time than ITL (137 seconds vs. 1 second) and showed poorer CVaR values, highlighting the challenges of gradient optimization in non-convex settings. ITL’s efficiency is also evident in Randomworld, with training times of 0.70 seconds compared to MCE’s 35 seconds (See Table 4). Our methods consistently converge to optimal performance as coverage increases, demonstrating consitency with expert demonstrations. MCE’s tendency to get trapped in local optima underscores the efficiency and reliability of ITL and BITL in complex environments.
The policies from our method stay in the support of the expert’s actions in every visited state, by imposing hard constraints leading to better worst case scenarios. Our methodology enforces hard constraints, ensuring policies adhere to expert actions in every visited state, thus recovering an $\epsilon$-optimal action for each state and minimizing constraint violations. Unlike MLE, PS, and MCE, which occasionally breach constraints (See Table 4 and Table 1 ) , ITL and BITL maintain strict compliance, enhancing reliability. CVaR results in Table 4 show ITL and BITL provide better worst-case outcomes by enforcing $\epsilon$-ball property on the estimated dynamics (See Definition 3), preserving policy robustness, even in challenging scenarios.
Our method integrates efficiently with Bayesian inference, useful for settings requiring more exploration and for predicting when we can transfer well Our integration with Bayesian inference enables us to infer calibrated uncertainties effectively, leveraging expert knowledge to enhance performance, particularly in environments like Gridworld where exploration avoid repeated suboptimal decisions such as repeatedly colliding with a wall due to poor action choices. This problem is less pronounced in Randomworld, illustrating the adaptability of our approach to different dynamics. Furthermore, the Bayesian framework allows us to predict with greater accuracy which tasks will likely yield successful outcomes based on the learned distributions. Looking at the Regrets plots (right column of Figure 2), we observe a clear correlations between the Bayesian Regret values and how well we perform on a task. The plot clearly show higher regret for the transfer task compared to the standard task which makes sense. In table 1, we show how similar reasoning can be applied to a real-life dataset where we see that we clearly perform better on task which lower regret.
Our method can be used into a continous real-life healthcare settings and provides insights through counterfactual reasoning. In the healthcare environment, we discretize the state space based on the following feature values: Oxygen ratio (O2), Blood pressure (BP), Creatinine (Crea), and Glasgow Coma Scale (GCS). Details on the discretization rules are provided in Table 2 of the Appendix. After prescribing intravenous treatment under the current state (O2 = 1, BP = 1, GCS = 1, Crea = 2), we examine the three most likely subsequent states for the MLE, ITL, and MCE methods. Intravenous treatment is expected to improve blood pressure and creatinine levels. Since clinicians never prescribed intravenous treatment in this state, MLE results are uniform across all states and omitted from the plots. ITL shows that the likelihood of the next states is distributed among states that either improve blood pressure, creatinine, or remain the same. In contrast, MCE assigns all its likelihood to a single state, which is unrealistic given the complexities of human physiology and the lack of observed transitions. This example highlights how the MCE framework can produce unrealistic results in complex real-life settings. It is important to note that for ITL, the probability of the next states in Figure 3 is close to uniform quantitatively. This is due to the fact that we are using a uniform prior and have not observed any data for that state. However, our method can rank and hence quantify the uncertainty over the next states in a more informed way than MLE (which is exactly uniform) or MCE, which reduces its uncertainty to 0 even when no data exists for that particular state.
Discussion
Summary We tackled the challenge of estimating transition dynamics $T^{*}$ from near-optimal expert trajectories in offline model-based reinforcement learning. Our approach, Inverse Transition Learning (ITL), leverages the limited coverage of expert trajectories and the near-optimality of the expert to estimate $T^{*}$. We introduced a novel constraint-based method and integrated these constraints within a Bayesian framework to learn a posterior distribution over the dynamics. To our knowledge, this is the first method to combine posterior estimation of dynamics with expert demonstrations. Our approach significantly enhances decision-making in both synthetic environments and real-world healthcare scenarios, such as ICU patient management for hypotension. It not only improves decision quality but also offers insights into the likelihood of successful task transfer, showcasing the robustness and adaptability of our method.
Future Work While our method provides a robust approach for estimating environmental dynamics using expert demonstrations, it is currently limited to discrete and fully observable state spaces. Future research could explore extending ITL to handle more complex environments, including those with high-dimensional, continuous, or partially observable state spaces. Additionally, combining ITL with Inverse Reinforcement Learning (IRL) to simultaneously learn rewards and dynamics represents another promising avenue for future work.
Appendix A Proofs
Theorem 1.
@@ Appendix A Proofs
Proof.
@@
$Q^{\pi^{*}(T^{*})}(s,a^{*};T)-Q^{\pi^{*}(T^{*})}(s,a^{\prime};T)\geq\epsilon\\ \implies Q^{\pi^{*}(T^{*})}(s,a^{*};T)\geq Q^{\pi^{*}(T^{*})}(s,a^{\prime};T),$
Thus, the transition dynamics $T$ will favor the same action $a^{*}$ as the true transition dynamics $T^{*}$ for each state $s$. So if we attempt policy improvement on $Q^{\pi^{*}(T^{*})}(s,.;T)$, we will end up with the same policy $\pi^{*}(T^{*})$, meaning $\pi^{*}(T)=\pi^{*}(T^{*})$ is the optimal policy for dynamics $T$. ∎
Lemma 1.
For any $\epsilon$, if $T$ satisfies constraints($\pi^{*}(T)$), then $T$ will satisfies the $\epsilon$-ball property.
Proof.
For each state $s$, Constraint 1($\pi^{*}(T)$) and Theorem 1 imply $\pi^{*}(T)(s)\in\epsilon(s;T^{*})$ because $Q^{*}(s,a;T)$ is greater than $Q^{*}(s,a^{\prime};T)+\epsilon$ for all $a\in\epsilon(s;T^{*})$ and $a^{\prime}\notin\epsilon(s;T^{*})$. This ensures that every invalid action is $\epsilon$-away from every valid action under the dynamics $T$. Additionally, Constraint 2($\pi^{*}(T)$) guarantees that $\epsilon(s;T^{*})=\epsilon(s;T)$ by ensuring that for each valid action $a\in\epsilon(s;T^{*})$, the $\max_{a^{\prime}}Q^{*}(s,a^{\prime};T)-Q^{*}(s,a;T)\leq\epsilon$. This completes the proof that the dynamics $T$ satisfy the $\epsilon$-ball property. ∎
Appendix B Non-Convex Feasible Region

### v2

@@ 6 EXPERIMENTAL SETUP
Environments Our evaluations span 11 synthetic environments and a real-life ICU setting. Specifically, we use a 25 states Gridworld with four actions, 10 different 15-states Randomworld with five actions each, and a healthcare scenario focusing on ICU patients with hypotension, utilizing the MIMIC-IV dataset (Johnson et al.,, 2020). In Gridworld, we generate 100 batches of data $\mathcal{D}$, each consisting of five episodes with 15 steps each. For each 10 of the Randomworld, 50 batches are generated, each containing three episodes of ten steps, due to their smaller size. We also evaluate our methods across varying ”Coverage %” levels in these environments, defined as the percentage of states observed within each batch data $\mathcal{D}$ and various $\epsilon$ values to see how our method compare to baselines for various degrees of sub-optimality. Detailed descriptions of both the synthetic and real-world environments are available in Appendix C and D respectively.
@@
In synthetic scenarios, we simulate a suboptimal expert, where approximately $40\%$ of states are stochastic-policy states (See definition 4), demonstrating the robustness of our approach to suboptimal expert behavior. We also include in the Appendix (Table 7, 8, Figure 8 and 9), results for $20\%$ and $0\%$ (fully optimal expert) of stochastic-policy states which correspond to lower $\epsilon$ values. Note that increased optimality from the expert (lower $\epsilon$ values) leads to less stochasticity (ie. less stochastic-policy states) amongst the actions selected in the batch data $\mathcal{D}$. Our method does not require the expert to be perfectly optimal and can adapt to various degrees of optimality through $\epsilon$. In the results below, we use $\gamma=0.95$ and smoothing parameter defined in Equation 3 $\delta=0.001$.
@@
Baselines. Our methods, Bayesian Inverse Transition Learning (BITL) and Inverse Transition Learning (ITL), are evaluated against $T^{MLE}$ (Maximum Likelihood Estimation, MLE), $T^{MCE}$ (Maximum Causal Entropy, MCE) (Herman et al.,, 2016), and a non-expert-informed posterior $P(T^{*}|\mathcal{D})$ (PS).
@@
[…]
Best: $\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}\mathcal{I}\!\left\{\arg\max_{a}\pi(a|s)=\arg\max_{a}\pi^{*}(a|s)\right\},$
$\epsilon\text{-ball: }$ $\frac{1}{|\mathcal{S}|}\sum_{s\in\mathcal{S}}\mathcal{I}\!\left\{\arg\max_{a}\pi(a|s)\in\epsilon(s;T^{*})\right\}.$
@@
$\frac{1}{|\widehat{P}(T|\mathcal{D})|^{2}}\sum_{T\in\widehat{P}(T|\mathcal{D})}\sum_{T^{\prime}\in\widehat{P}(T|\mathcal{D})}$
$\quad\times\mathbb{E}_{s_{0}\sim\mu_{0}}\!\left[\big|V^{\pi^{*}(T)}(s_{0};T)-V^{\pi^{*}(T^{\prime})}(s_{0};T)\big|\right]$
@@
Transfer Task Setting
In this paper, a transfer task refers to the evaluation of the learned dynamics and/or policy under a reward function different from the one used during training. This setting is particularly relevant in healthcare, where the transition dynamics, which describe the physiological response of the human body to treatments, remain largely consistent. However, the reward function, which captures the clinical objectives being optimized, can vary across hospitals, patient populations, and individual clinicians. By evaluating our method under different reward functions, we assess its robustness to variations in clinical goals, ensuring that the learned dynamics remain useful even when the optimization criteria shift. This aligns with real-world applications, where treatment strategies may change while the underlying patient physiology remains unchanged. To demonstrate that our learned dynamics can adapt to actions not seen in the demonstrations, we set up a transfer task where the optimal actions under this new reward function differ from those in the original task (See Figure 4, 5). This ensures the new task requires actions we didn’t observe in the initial demonstrations. This tests the model’s adaptability to changes in task conditions.
Computing the results. For each dynamics $T$, we compute its optimal policy $\pi^{*}(T)$ using Value Iteration (Sutton and Barto,, 2018), when needed to compute metrics in terms of policies. For posterior distributions $P(T|\mathcal{D})$ and $P_{\epsilon}(T|\mathcal{D})$, we determine the optimal policies of 5,000 sampled $T^{\prime}s$ and average them.
@@
Figure 1: Top row: Normalized Value vs. Coverage for Gridworld (left: Standard Task, middle: Transfer Task), Bottom row: Normalized Value vs. Coverage for Randomworlds (left: Standard Task, middle: Transfer Task). Rightmost plots: Normalized Value vs. Bayesian Regret of both Tasks (top: Gridworld, bottom: Randomworlds). Overall ITL and BITL outerperforms other baselines and BITL help predicting when we can transfer well.
Gridworld
Method MCE BITL ITL MLE PS
$\epsilon$ matching $0.65\pm 0.12$ $\mathbf{0.78\pm 0.08}$ $0.65\pm 0.12$ $0.37\pm 0.06$ $0.38\pm 0.02$
$\epsilon$ matching transfer $0.5\pm 0.11$ $\mathbf{0.64\pm 0.08}$ $0.51\pm 0.11$ $0.39\pm 0.07$ $0.39\pm 0.02$
Best matching $0.55\pm 0.11$ $\mathbf{0.64\pm 0.08}$ $0.56\pm 0.11$ $0.31\pm 0.06$ $0.29\pm 0.02$
Best matching transfer $0.45\pm 0.1$ $\mathbf{0.54\pm 0.08}$ $0.45\pm 0.11$ $0.34\pm 0.07$ $0.31\pm 0.02$
[…]
Total Variation $137.62\pm 4.39$ $159.36\pm 3.98$ $\mathbf{137.05\pm 4.32}$ $141.37\pm 3.8$ $160.05\pm 4.43$
Value CVaR 5% $108.76\pm 26.54$ $109.04\pm 4.26$ $\mathbf{109.7\pm 9.01}$ $4.71\pm 0.47$ $2.35\pm 1.35$
Nbr constraints violated $3.06\pm 3.68$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$ $23.13\pm 6.59$ $16.37\pm 3.21$
Randomworlds
Method MCE BITL ITL MLE PS
$\epsilon$ matching $0.54\pm 0.12$ $0.75\pm 0.12$ $\mathbf{0.76\pm 0.13}$ $0.43\pm 0.11$ $0.3\pm 0.05$
$\epsilon$ matching transfer $0.46\pm 0.12$ $0.38\pm 0.1$ $\mathbf{0.5\pm 0.13}$ $0.46\pm 0.12$ $0.31\pm 0.05$
Best matching $0.38\pm 0.13$ $0.57\pm 0.12$ $\mathbf{0.58\pm 0.15}$ $0.29\pm 0.11$ $0.19\pm 0.05$
Best matching transfer $0.34\pm 0.13$ $0.26\pm 0.09$ $\mathbf{0.37\pm 0.15}$ $0.34\pm 0.13$ $0.2\pm 0.05$
Time $36.61\pm 26.36$ $5561\pm 223.43$ $65\pm 0.39 0.0$$0\mathbf$$\pm 0.0$} $2.61\pm 0.36$
Total Variation $111.08\pm 2.42$ $123.3\pm 4.06$ $\mathbf{102.02\pm 4.84}$ $111.07\pm 2.3$ $127.02\pm 2.66$
Value CVaR 5% $459.71\pm 23.48$ $364.34\pm 34.06$ $\mathbf{-366.43\pm 16.03}$ $481.98\pm 23.31$ $434.24\pm 17.22$
Nbr constraints violated $11.81\pm 6.47$ $\mathbf{0.0\pm 0.0}$ $\mathbf{0.0\pm 0.0}$ $7.23\pm 6.75 11$$1.66\pm 3.12$
Table 1: Gridworld and Randomworld results. BITL performs best in Gridworld, ITL in Randomworld.
7 RESULTS
ITL consistently outperforms baseline methods across all metrics and coverage settings. Table 1 (synthetic environments), Table 2 (ICU dataset), and Figure 1 show that ITL and BITL outperform baselines across metrics and coverage levels in both synthetic and real healthcare settings. Additional results for other $\epsilon$ values are provided in the appendix (Tables 7, 8, Figures 8, 9). In Randomworlds, BITL and ITL clearly surpass MCE, which itself outperforms MLE and PS. In Gridworld, ITL slightly outperforms MCE on average but achieves substantially better worst-case performance, underscoring the value of enforcing hard constraints and the risk of MCE getting stuck in poor local optima due to gradient-based optimization. In the ICU hypotension management task, ITL and BITL significantly outperform all baselines (Table 2), demonstrating the robust applicability of our approach to high-stakes, data-limited clinical settings. ITL also trains substantially faster. In the standard task, ITL and BITL are the only methods that converge to optimal performance as coverage increases (Figure 1, left column), achieving higher values without constraint violations—thanks to our hard-constraint design versus MCE’s soft constraints.
ITL enables reliable, fast, and data-efficient dynamics estimation. In Gridworld, MCE required far more training time than ITL (137s vs. 1s) and produced poorer CVaR values, reflecting the difficulties of gradient optimization in non-convex settings. ITL’s efficiency also holds in Randomworld, with training times of 0.7s versus 35s for MCE (Table 6). Our methods consistently converge to optimal performance as coverage increases, aligning with expert demonstrations. MCE’s tendency to get stuck in local optima further underscores the efficiency and reliability of ITL and BITL in complex environments.
Our policies stay within the support of expert actions in every visited state, with hard constraints yielding better worst-case outcomes. By enforcing hard constraints, our method ensures policies adhere to expert actions in every visited state, recovering an $\epsilon$-optimal action and minimizing violations. In contrast, MLE, PS, and MCE occasionally breach constraints (Tables 6, 2), while ITL and BITL maintain strict compliance, enhancing reliability. CVaR results (Table 6) show ITL and BITL achieve stronger worst-case performance by enforcing the $\epsilon$-ball property on estimated dynamics (Definition 3), preserving robustness even in challenging real-world scenarios such as ICU hypotension management.
Our method integrates efficiently with Bayesian inference, enabling exploration and predicting transfer success. By inferring calibrated uncertainties, BITL leverages expert knowledge to enhance performance in environments like Gridworld, where it avoids repeated suboptimal decisions (e.g., wall collisions) and achieves the best overall results. This shows that modeling uncertainty over dynamics while incorporating demonstrations provides a clear advantage over point-estimate methods. The issue is less pronounced in Randomworld and the ICU, highlighting the adaptability of our approach across different dynamics. The Bayesian framework also enables more accurate predictions of which tasks will yield successful outcomes based on the learned distributions. Regret plots (Figure 1, right column) show a strong correlation between Bayesian regret values and task performance, with higher regret in transfer tasks than in standard tasks, as expected. Table 2 further illustrates this in the ICU dataset, where tasks with lower regret align with better performance.
Our method applies to continuous healthcare settings and provides insights through counterfactual reasoning. In the ICU environment, we discretize states using Oxygen ratio (O2), Blood pressure (BP), Creatinine (Crea), and Glasgow Coma Scale (GCS) (Table 4 in Appendix). After prescribing IV treatment in state (O2=1, BP=1, GCS=1, Crea=2), we examine the three most likely next states under MLE, ITL, and MCE. Since clinicians never prescribed IV here, MLE gives uniform predictions and is omitted. ITL spreads probability across states that improve or maintain BP and creatinine, while MCE collapses all probability on a single state—unrealistic given physiological complexity and missing data. This illustrates how MCE can yield implausible results in real-world settings. ITL’s next-state probabilities in Figure 2 appear near-uniform due to the prior and absence of data, yet ITL still ranks states and quantifies uncertainty more informatively than MLE (exactly uniform) or MCE (which reduces uncertainty to zero).
Figure 2: Most likely next three states after IV in state O2=1, BP=1, GCS=1, Crea=2. MLE omitted (uniform). ITL spreads mass over improving/maintaining states; MCE collapses onto one state.
Method MLE ITL MCE BITL PS
Standard Task
Best match. 0.33 0.51 0.31 0.49 0.31
$\epsilon$ match. 0.52 1.0 0.56 1.0 0.51
Constraints 61 0 47 0 58
Time (s) - 2.23 118 - -
Bayes Regret - - - 2.49 10
Transfer Task
Best match. 0.34 0.52 0.32 0.47 0.34
$\epsilon$ match. 0.58 0.97 0.68 0.90 0.58
Bayes Regret - - - 2.80 5.40
Table 2: ICU results ($\epsilon=5$). ITL beats all point-estimates and BITL surpasses all Bayesian-estimates.
8 DISCUSSION
Summary. We addressed the challenge of estimating transition dynamics $T^{*}$ from near-optimal expert trajectories in offline model-based RL. Our approach, Inverse Transition Learning (ITL), leverages limited expert coverage and near-optimality to estimate $T^{*}$. We introduced a novel constraint-based method and integrated it within a Bayesian framework to learn a posterior over the dynamics. To our knowledge, this is the first method to combine posterior estimation of dynamics with expert demonstrations. ITL improves decision-making in both synthetic environments and real-world healthcare, such as ICU hypotension management. It not only enhances decision quality but also provides insights into task transfer.
Future Work. Although ITL provides a robust approach for estimating dynamics from demonstrations, it is currently limited to discrete, fully observable state spaces. Future work could extend ITL to high-dimensional, continuous, or partially observable settings, along with deeper theoretical analysis.
Acknowledgments
We thank the fantastic AISTATS reviewers for their constructive comments and suggestions. This material is based upon work supported by the National Science Foundation under Grant No. IIS-2007076 as well as the Cooperative AI Foundation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation nor the Cooperative AI Foundation.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Not Applicable]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [No]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Not Applicable]
    2. (b)
      Complete proofs of all theoretical results. [Not Applicable]
    3. (c)
      Clear explanations of any assumptions. [Not Applicable]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [No]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator if your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Bayesian Inverse Transition Learning:
Learning Dynamics from Near-Optimal Trajectories
Supplementary Materials
Appendix A Non-Convex Feasible Region